For the reaction: FeCl 3(aq) + 3NaOH(aq) → Fe(OH) 3(s) + 3NaCl(aq) What mass of Fe(OH) 3 can be produced from 135 mL of 0.04 mol/L NaOH? 


1
Expert's answer
2021-01-18T01:46:44-0500

Solution:

The balanced chemical equation:

FeCl3(aq) + 3NaOH(aq) → Fe(OH)3(s) + 3NaCl(aq)

According to the equation above: n(NaOH)/3 = n(Fe(OH)3)


Moles of NaOH = Molarity of NaOH × Solution volume

n(NaOH) = Moles of NaOH = 0.04 mol/L × 0.135 L = 0.0054 mol


Moles of Fe(OH)3 = n(Fe(OH)3) = n(NaOH)/3

Moles of Fe(OH)3 = 0.0054 mol / 3 = 0.0018 mol


Moles of Fe(OH)3 = Mass of Fe(OH)3 / Molar mass of Fe(OH)3

The molar mass of Fe(OH)3 is 106.867 g mol-1.

Hence,

Mass of Fe(OH)3 = Moles of Fe(OH)3 × Molar mass of Fe(OH)3

Mass of Fe(OH)3 = 0.0018 mol × 106.867 g mol-1 = 0.19236 g = 0.19 g

Mass of Fe(OH)3 = 0.19 g


Answer: 0.19 grams of Fe(OH)3 can be produced.